# HyWA: Hypernetwork Weight Adapting Personalized Voice Activity Detection

### v1

@@ 3.4 Results
Table LABEL:tab:Average_precision presents a comparison between our proposed approach HyWA, and four widely-used PVAD configurations for speaker conditioning: (1) Concatenation, (2) Multiplication, (3) Addition, and (4) Feature-wise linear modulation (FiLM) [5, 15, 16], see Figure 3.
We consider three experimental scenarios. The first scenario, named “Clean”, corresponds to a noise-free environment, where the dataset contains only clean speech signals. This setup allows us to examine the inherent capacity of each method without the confounding effect of noise. The second scenario (Seen Noise) reflects a condition where the test set includes added noise during training. Finally, the (Unseen Noise) scenario evaluates generalization to novel acoustic conditions by incorporating noise types that were not observed during training. For noisy cases, we report results averaged over different SNR levels, thereby providing a comprehensive measure of robustness.
Table LABEL:tab:Average_precision confirms that HyWA improves mAP and AP in all four baselines across all scenarios: clean, seen noise, and unseen noise. Under “seen noise”, and “unseen noise” our approach continues to outperform other competing methods, which assures the robustness and the signal extraction capacity of our PVAD. Integrating a hypernetwork significantly enhances the personalization of VAD systems, ensuring accurate target-speaker detection even under challenging acoustic environments.
@@ 4 Conclusion
We introduced a hypernetwork-based speaker conditioning method to personalize voice activity detection. Our method performs user personalization through weight adaptation, without altering the underlying VAD architecture. By selectively modifying the weights of a few layers using speaker embeddings, HyWA achieves effective personalization while maintaining architectural simplicity and deployment flexibility. The ability to reuse a single base VAD model across different speakers without retraining or redesigning VAD offers a simple and scalable solution for real-world applications while improving model performance.

### v2

@@ 3.4 Results
Table 1 compares our proposed approach, HyWA, against four commonly used speaker-conditioning schemes for PVAD: (1) Concatenation, (2) Multiplication, (3) Addition, and (4) feature-wise linear modulation (FiLM) [personal_vad_2, SSL-pvad-baseline-journalVersion, comparative-analysis-personalized-voice] (see Figure 3). For FiLM, note that we evaluated conditioning at multiple insertion sites, including pre- and post-LSTM block. We report the pre-LSTM FiLM configuration because it achieves 1.5% higher mAP than post-LSTM in our setting, consistent with observations in [SSL-pvad-baseline-journalVersion].
We consider three experimental scenarios. The first, ``Clean'', corresponds to a noise-free environment, where the dataset contains only clean speech signals. This setup allows us to examine the inherent capacity of each method without the confounding effect of noise. The second, ``Seen Noise'', reflects a condition where the test set includes noise types observed during training. Finally, the ``Unseen Noise'' scenario evaluates generalization to novel acoustic conditions by incorporating noise types that were not observed during training. For both noisy scenarios, we generate test utterances using the same SNR range and sampling procedure as in training, namely 5-5 to $20$ dB in $5$ dB increments. We report results averaged over SNR levels in Table 1, thereby providing a comprehensive measure of robustness.
Table 1 indicates that HyWA improves AP and mAP over all four baselines across all scenarios. In the ``Clean'' condition, our model attains the best AP and mAP, demonstrating strong target-speaker discrimination in ideal settings. Under both ``Seen Noise'' and ``Unseen Noise'', our approach continues to outperform other competing methods, suggesting improved robustness and better target-speaker extraction. Overall, these results indicate that integrating a hypernetwork significantly enhances the personalization of VAD systems and enables more accurate target-speaker detection even under challenging acoustic environments.
@@ 4 Conclusion
We introduced HyWA, a new speaker-conditioning method to personalize voice activity detection. This view shifts personalization from modifying inputs or activations of a VAD, towards generating user-specific weights. HyWA performs personalization with a one-time forward pass of the hypernetwork given an enrollment utterance. The hypernetwork is trained to maximize PVAD performance for a VAD backbone. This allows HyWA to learn a mapping from the enrollment-derived speaker embeddings to speaker-specific subsets of VAD weights. Comparison with conditioning baselines shows that HyWA delivers consistent PVAD improvements through higher mean average precision. This perspective allows to reuse a single base VAD model across different speakers, without retraining or redesigning VAD. Overall, HyWA provides a simple and effective technique to personalize a VAD model. Weight-generation-based conditioning is a promising direction for future personalized speech detection systems.